# On automatic boundedness of operators between ordered and topological vector spaces

### 11 Mar 2026 v2

@@ 1. $i)$
    If $Y$ is a Banach space, then an operator $T:X\to Y$ is continuous whenever it takes order intervals onto relatively weak compact sets ([14, Corollary 2.9]).
@@
    If the cone of $X$ is also normal and $Y$ is an NS, then $\text{\bf L}_{rnc}(X,Y)=\text{\bf L}_{onb}(X,Y)=\text{\bf L}_{b}(X,Y)$ ([14, Corollary 2.12]).
@@ 3. $iii)$
    If the cone of $X$ is also normal and $Y$ is an ONS with a generating normal cone, then ${\cal L}_{rc}(X,Y)\subseteq{\cal L}_{b}(X,Y)$ ([14, Corollary 2.14]).
@@ 4. $iv)$
    If $Y$ is an OBS with a closed generating cone and $X^{\sim}\neq\emptyset$, then every ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ is uniformly bounded iff $Y_{+}$ is normal ([14, Proposition 2.15]).
Let $X$ be an OBS with a normal cone and $Y=(Y,\tau)$ be a Banach space with the weak topology.
@@ 1. $v)$
    ${\cal L}_{o{\tau}c}^{\sigma}(X,Y)\subseteq{\cal L}_{onb}(X,Y)$ ([14, Proposition 2.16]).
@@ 2. $vi)$
    If the cone of $X$ is also closed and generating, then every order-to-topology $\sigma$-continuous operator $T:X\to Y$ is bounded ([14, Corollary 2.18]).
The following theorem generalizes Proposition 2.5 to the setting of ordered Fréchet spaces. The idea of its proof below is motivated by the proof of [14, Theorem 2.5]. Recall that a Fréchet space $({\cal F},\tau)$ is an LCS, whose locally convex topology $\tau$ is metrizable by a complete metric.
@@
We have $\text{\bf L}_{o{\tau}b}({\cal F},Y)\subseteq\text{\bf L}_{b}({\cal F},Y)$, whenever $({\cal F},\tau)$ is an ordered Fréchet space with a $\tau$-closed generating cone and $(Y,\xi)$ is a TVS.
@@
Every order-to-topology bounded operator from an ordered Fréchet space with a closed generating cone to a TVS is topologically bounded.
Since $\text{\bf L}_{b}(X,Y)\subseteq\text{\bf L}_{o{\tau}b}(X,Y)$ whenever $X$ is an ordered TVS with a normal cone and $Y$ is a TVS, the following result follows from Proposition 2.3 and Theorem 2.7.
@@
[…]
  1. $i)$
    $\text{\bf L}_{r{\tau}c}({\cal F},Y)=\text{\bf L}_{r{\tau}b}({\cal F},Y)=\text{\bf L}_{b}({\cal F},Y)$.
  2. ii)ii)
    ${\cal L}_{r{\tau}c}({\cal F},Y)={\cal L}_{o{\tau}b}({\cal F},Y)={\cal L}_{b}({\cal F},Y)$.
By [16, Theorem 2.4], ${\cal L}_{rc}(X,Y)={\cal L}_{ob}(X,Y)$ whenever OVSs $X$ and $Y$ both have generating cones. Keeping in mind that ${\cal L}_{ob}(X,Y)\subseteq{\cal L}_{o{\tau}b}(X,Y)$ if $Y$ is an ordered TVS with a normal cone, we obtain the following consequence of Corollary 2.9 ii)ii) providing conditions for automatic topological boundedness of ru-continuous operators.
Corollary 2.10.
[…]
3 On boundedness of Levi and Lebesgue operators between ordered TVSs
We use the following versions of $\tau$-Levi and $\tau$-Lebesgue operators (see [5, 13, 20, 22] in the vector lattice-setting).
@@ 3. $d)$
    $(\sigma$-))order-to-weak continuous if $Tx_{\alpha}\operatorname{\xrightarrow[]{\sigma(Y,Y^{\prime})}}0$ for every ((sequence)) net $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$. The set of such operators is denoted by ${\cal L}_{owc}(X,Y)$ (resp., ${\cal L}^{\sigma}_{owc}(X,Y)$) (cf. Definition 2.1d)d)).
We prove that, under a rather mild condition, quasi $\sigma$-Levi operators are topologically bounded. This result generalizes [13, Theorem 3.1].
@@
Let $T$ be a quasi $\sigma$-Levi operator from an ordered Fréchet space $({\cal F},\tau)$ with a $\tau$-closed generating cone to an ordered TVS $(Y,\xi)$ with a normal cone. Then $T$ is topologically bounded.
@@
Recall that a positive operator $T$ from a vector lattice to a normed lattice is $\sigma-Lebesgue$ iff $T$ is $\sigma-order-to-norm continuous$ [5, Lemma 2.1]. The same result holds also in the following more general setting.
@@
$T\in{\cal L}_{Leb}(X,Y)\Longleftrightarrow T\in{\cal L}_{o{\tau}c}(X,Y)\ \ \ and\ \ \ T\in{\cal L}^{\sigma}_{Leb}(X,Y)\Longleftrightarrow T\in{\cal L}^{\sigma}_{o{\tau}c}(X,Y).$
@@
As the $\sigma$-Lebesgue case is similar, we consider only the first equivalence. Since obviously ${\cal L}_{o{\tau}c}(X,Y)\subseteq{\cal L}_{Leb}(X,Y)$, it suffices to show that $T\in{\cal L}_{Leb}(X,Y)$ implies $T\in{\cal L}_{o{\tau}c}(X,Y)$.
On the way to a contradiction, suppose $T\in{\cal L}_{Leb}(X,Y)\setminus{\cal L}_{o{\tau}c}(X,Y)$. Then, there exist a net $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$ and a neighborhood $U\in\tau(0)$ such that, for every $\alpha$, there exists $\alpha^{\prime}\geq\alpha$ with $Tx_{\alpha^{\prime}}\notin U$. Since $Y_{+}$ is normal, we may suppose that $U$ is circled and full by [4, Corollary 2.21].
Pick a net $g_{\beta}\downarrow 0$ in $X$ such that, for each $\beta$, there exists $\alpha_{\beta}$ with $x_{\alpha}\in[-g_{\beta},g_{\beta}]$ for $\alpha\geq\alpha_{\beta}$. As $T$ is Lebesgue, $Tg_{\beta}\operatorname{\xrightarrow[]{\tau}}0$. Since $T\geq 0$, $Tg_{\beta}\downarrow\geq 0$. Suppose $Tg_{\beta}\geq f\geq 0$ for some $f\in Y$. Since $\tau$ is a Hausdorff topology, it follows from [4, Theorem 2.23] that $f=0$, and hence $Tg_{\beta}\downarrow 0$. Since $T\geq 0$, we obtain $Tx_{\alpha}\in[-Tg_{\beta},Tg_{\beta}]$ for $\alpha\geq\alpha_{\beta}$. Since $Tg_{\beta}\operatorname{\xrightarrow[]{\tau}}0$, there exists $\beta_{U}$ such that $Tg_{\beta}\in U$ for all $\beta\geq\beta_{U}$. Since $U$ is circled and full, $Tg_{\beta_{U}},Tg_{\beta_{U}}]\subseteq U$ for all $\beta\geq\beta_{U}$, and hence $Tx_{\alpha}\in U$ for all $\alpha\geq\alpha_{\beta_{U}}$. It is absurd because $Tx_{\alpha^{\prime}_{\beta_{U}}}\notin U$. Thus, $T\in{\cal L}_{o{\tau}c}(X,Y)$. ∎
@@
By [5, Proposition 2.2], a positive order continuous operator $T$ from a vector lattice to a normed lattice is Lebesgue iff $T$ is w-Lebesgue. We extend this fact to OVS-setting as follows.
@@
Question 3.7.
Are the conclusions of Corollaries 3.4 and 3.6 hold true for arbitrary (not necessarily positive) operators?
Till the end of the section, we restrict ourselves to operators with a normed space in the co-domain. The following lemma is a direct consequence of Definition 3.1.
@@
${\cal L}_{o{\tau}c}(X,Y)\subseteq{\cal L}_{owc}(X,Y),\ \ {\cal L}^{\sigma}_{o{\tau}c}(X,Y)\subseteq{\cal L}^{\sigma}_{owc}(X,Y),$
${\cal L}_{o{\tau}c}(X,Y)\subseteq{\cal L}_{Leb}(X,Y)\bigcap{\cal L}^{\sigma}_{o{\tau}c}(X,Y)\subseteq{\cal L}_{Leb}(X,Y)\bigcup{\cal L}^{\sigma}_{o{\tau}c}(X,Y)\subseteq{\cal L}^{\sigma}_{Leb}(X,Y),\ \ \text{\rm and}$
@@
Let $X$ be an OBS with a closed generating normal cone and $Y$ be an NS. Then ${\cal L}_{onb}(X,Y)={\rm L}(X,Y)$.
@@
The inclusion ${\cal L}_{onb}(X,Y)\subseteq{\rm L}(X,Y)$ follows from [16, Corollary 2.9]. The reverse inclusion hods due to normality of $X$. ∎
By [16, Theorem 2.8], every order-to-norm bounded operator from an OBS with closed generating cone to an NS is bounded. Accordingly to [16, Theorem 2.4], each order bounded operator from an OVS to an OVS with generating cones is ru-continuous. Combining these two results and keeping in mind that order intervals in an OBS with normal cone are bounded, we obtain the following automatic boundedness conditions of order bounded operators:
@@
    ${\cal L}_{ob}(X,Y)\subseteq{\rm L}(X,Y)$ whenever $X$ and $Y$ are OBSs with generating cones such that $X_{+}$ is closed and $Y_{+}$ is normal.
Another case of automatic boundedness of operators concerns the order continuity. Namely, as every order continuous operator from an Archimedean OVS with generating cone to an OVS is order bounded [16, Corollary 2.2], we have:
@@
    ${\cal L}_{oc}(X,Y)\subseteq{\rm L}(X,Y)$ whenever $X$ and $Y$ are OBSs with generating cones such that $X_{+}$ is closed and $Y_{+}$ is normal.
Condition $X_{+}-X_{+}=X$ is essential in both (*) and (**) since each $T:X\to X$ is order bounded and order continuous whenever $X_{+}=\{0\}$. An inspection of [16, Example 2.12 b)] tells us that each infinite dimensional Banach space $X$ has a non-closed generating Archimedean cone so that every operator $T:X\to X$ is order continuous, and hence order bounded.
@@
Combining Lemmas 3.8, 3.9 and 3.10, we obtain.
@@
Let $X$ be an OBS with a closed generating normal cone and $Y$ be an NS. Then ${\cal L}^{\sigma}_{wLeb}(E,Y)\subseteq{\rm L}(E,Y)$. In particular, ${\cal L}^{\sigma}_{ows}(E,Y)\subseteq{\rm L}(E,Y)$.
@@
(cf. [5, Proposition 3.2], [23, Lemma 2.3]) Each $\sigma$-w-Lebesgue, and hence each $\sigma$-order-to-weak continuous, operator from a Banach lattice to an NS is bounded.
@@
Let $X$ be an OBS with a closed generating normal cone and $Y$ be an NS. Then ${\cal L}_{Leb}(X,Y)$, ${\cal L}^{\sigma}_{Leb}(X,Y)$, ${\cal L}_{wLeb}(X,Y)$, ${\cal L}^{\sigma}_{wLeb}(X,Y)$, ${\cal L}_{onc}(X,Y)$, ${\cal L}^{\sigma}_{onc}(X,Y)$, ${\cal L}_{owc}(X,Y)$, and ${\cal L}^{\sigma}_{owc}(X,Y)$ are norm closed subspaces of ${\rm L}(X,Y)$.
Positive order-to-norm continuous order-to-norm bounded operators between normed lattices are not necessary bounded (see [5, Remark 1.1. c)]). From the other hand, even a positive $\sigma$-Lebesgue rank one contraction on an AM-space need not to be w-Lebesgue [5, Example 3.1)].

### 16 Apr 2026 v3

@@ 1. $i)$
    If $Y$ is a Banach space, then an operator $T:X\to Y$ is continuous whenever it takes order intervals onto relatively weak compact sets [12, Corollary 2.9].
@@
    If $X_{+}$ is normal and $Y$ is a NS, then $\text{\bf L}_{rnc}(X,Y)=\text{\bf L}_{onb}(X,Y)=\text{\bf L}_{b}(X,Y)$ [12, Corollary 2.12].
@@ 3. $iii)$
    If $X_{+}$ is normal and $Y$ is an ONS with a generating normal cone, then ${\cal L}_{rc}(X,Y)\subseteq{\cal L}_{b}(X,Y)$ [12, Corollary 2.14].
@@ 4. $iv)$
    If $Y$ is an OBS with a closed generating cone and $X^{\sim}\neq\emptyset$, where $X^{\sim}$ is an order dual of $X$, then every ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ is uniformly bounded if and only if $Y_{+}$ is normal [12, Proposition 2.15].
@@ 5. $v)$
    If $X_{+}$ is normal and $Y=(Y,\tau)$ is a Banach space with the weak topology, then ${\cal L}_{otc}^{\sigma}(X,Y)\subseteq{\cal L}_{onb}(X,Y)$ [12, Proposition 2.16].
@@ 6. $vi)$
    If $X_{+}$ is closed generating and normal and $Y=(Y,\tau)$ is a Banach space with the weak topology, then every order-to-topology $\sigma$-continuous operator $T:X\to Y$ is topologically bounded [12, Corollary 2.18].
[…]
@@
$\text{\bf L}_{otb}({\cal F},Y)\subseteq\text{\bf L}_{c}({\cal F},Y)$, whenever $({\cal F},\tau)$ is an ordered Fréchet space with a $\tau$-closed generating cone and $(Y,\xi)$ is a topological vector space.
@@
Suppose, on the contrary, ${\cal T}\in\text{\bf L}_{otb}({\cal F},Y)\setminus\text{\bf L}_{c}({\cal F},Y)$. For each $n\in\mathbb{N}$, denote
$V_{n}:=B_{\cal F}(0,2^{-2n})\cap{\cal F}_{+}-B_{\cal F}(0,2^{-2n})\cap{\cal F}_{+}.$
Clearly, $\{B_{\cal F}(0,2^{-2n}):n\in\mathbb{N}\}$ is a $\tau$-neighborhood base of zero consisting of $\tau$-closed sets and $B_{\cal F}(0,2^{-2(n+1)})+B_{\cal F}(0,2^{-2(n+1)})\subseteq B_{\cal F}(0,2^{-2n})$ holds for each $n$. Furthermore, each $B_{\cal F}(0,2^{-2n})$ is circled. To see this, it suffices to show that $p(x,0)\leq 2^{-2n}$ implies $p(tx,0)\leq 2^{-2n}$ whenever $|t|\leq 1$. So, let $p(x,0)\leq 2^{-2n}$ and $|t|\leq 1$. Since $\frac{|t|a}{1+|t|a}\leq\frac{a}{1+a}$ for every $a\in\mathbb{R}_{+}$, we obtain $p(tx,0)=\sum\limits_{k=1}^{\infty}\frac{2^{-k}p_{k}(tx)}{1+p_{k}(tx)}=\sum\limits_{k=1}^{\infty}\frac{2^{-k}|t|p_{k}(x)}{1+|t|p_{k}(x)}\leq\sum\limits_{k=1}^{\infty}\frac{2^{-k}p_{k}(x)}{1+p_{k}(x)}=p(x,0)\leq 2^{-2n}$. By the Ando theorem [4, Theorem 2.10], $\{V_{n}:n\in\mathbb{N}\}$ is a $\tau$-neighborhood base of zero consisting of (not necessarily open) circled subsets satisfying $V_{n+1}+V_{n+1}\subseteq V_{n}$ for all $n$.
[…]
@@
Every order-to-topology bounded operator from an ordered Fréchet space with a closed generating cone to a topological vector space is continuous.
Keeping in mind that ${\cal L}_{ob}(X,Y)\subseteq{\cal L}_{otb}(X,Y)$ if $Y$ is a TVS with a normal cone [4, Lemma 2.22], we obtain the following conditions for automatic continuity of order bounded operators.
@@
Every order bounded operator from an ordered Fréchet space with a closed generating cone to an ordered locally convex space with a normal cone is continuous.
3 On continuity of the Levi and the Lebesgue operators
We shall use the following versions of $\tau$-Levi and $\tau$-Lebesgue operators (see also [5, 11, 19, 21] in the vector lattice-setting).
@@ 3. $d)$
    $(\sigma$-))order-to-weak continuous if $Tx_{\alpha}\operatorname{\xrightarrow[]{\sigma(Y,Y^{\prime})}}0$ for every ((sequence)) net $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$. The set of such operators is denoted by ${\cal L}_{owc}(X,Y)$ (resp., ${\cal L}^{\sigma}_{owc}(X,Y)$) (cf., Definition 2.1d)d)).
We prove that, under rather mild conditions, quasi $\sigma$-Levi operators are continuous. This result generalizes [11, Theorem 3.1].
@@
Every quasi $\sigma$-Levi operator from a Banach lattice to a normed lattice is continuous.
@@
Let $T:E\to F$ be a quasi $\sigma$-Levi operator from a BL $E$ to an NL $F$. On the way to a contradiction, suppose $T$ is not continuous. Then, there exists a sequence $(x_{n})$ in $B_{X}$ such that $\|Tx_{n}\|\to\infty$. Without loss of generality, suppose $(x_{n})\subset(B_{X})^{+}$ and $\|Tx_{n}\|\geq n2^{n}$ for every $n$. Set $u=\sum\limits_{k=1}^{\infty}2^{-k}x_{k}$ and $u_{n}=\sum\limits_{k=1}^{n}2^{-k}x_{k}$ for each $n\in\mathbb{N}$. Then $0\leq u_{n}\uparrow$ and $\|u_{n}\|\leq 1$ for all $n$. Since $T\in\text{\rm L}_{\text{\rm qLevi}}(E,F)$, $(Tu_{n})$ is an o-Cauchy sequence in $F$, and hence the sequence $(Tu_{n})$ is order bounded. Thus, $(Tu_{n})$ is also norm bounded. It is absurd because $\|Tu_{n+1}-Tu_{n}\|=\|T2^{-(n+1)}x_{n+1}\|\geq n+1$ for all $n$. ∎
Theorem 3.3.
Every quasi $\sigma$-Levi operator from an ordered Fréchet space with a closed generating cone to an ordered normed space with a normal cone is continuous.
Proof.
Let $T:{\cal F}\to Y$ be a quasi $\sigma$-Levi operator from an ordered Fréchet space $({\cal F},\tau)$ with a $\tau$-closed generating cone to an ONS $Y$ with a normal cone. Toward a contradiction, suppose that $T$ is not continuous.
@@
Letting $x_{n}:=2^{n}z_{n}$, denote $u_{i}=\sum\limits_{k=1}^{i}x_{k}$ for each $i\in\mathbb{N}$. Clearly, $(u_{i})$ is an increasing sequence. Take an $l\in\mathbb{N}$. It is straightforward to see that, for each $m\in\mathbb{N}$,
$\sum\limits_{k=l+1}^{m+l+1}x_{k}=\sum\limits_{k=l+1}^{m+l+1}2^{k}z_{k}\in\sum\limits_{k=l+1}^{m+l+1}2^{k}B_{\cal F}(0,2^{-2k})\subseteq\sum\limits_{k=0}^{m}2^{-k}B_{\cal F}(0,2^{-(l+1)})\subseteq B_{\cal F}(0,2^{-l}).$
Find an $N>0$ satisfying $\sum\limits_{k=1}^{j}x_{k}\in NB_{\cal F}(0,2^{-l})$ for j=1,,lj=1,...,l. Then, for every $i\in\mathbb{N}$,
$u_{i}=\sum\limits_{k=1}^{i}x_{k}=\sum\limits_{k=1}^{\min(i,l)}x_{k}+\sum\limits_{k=\min(i,l)+1}^{i}x_{k}\in(N+1)B_{\cal F}(0,2^{-l}).$
As $\{B_{\cal F}(0,2^{-l})\}_{l\in\mathbb{N}}$ is a $\tau$-neighborhood base of zero, the sequence $(u_{i})$ is $\tau$-bounded.
It follows from $T\in{\cal L}^{\sigma}_{\text{\rm qLevi}}({\cal F},Y)$ that $(Tu_{i})$ is o-Cauchy in $Y$, and hence is order bounded. Since $Y_{+}$ is normal, the sequence $(Tu_{i})$ is also norm bounded. It is absurd because
$\|Tu_{i+1}-Tu_{i}\|=\|Tx_{i+1}\|=\|T(2^{i+1}z_{i+1})\|\geq 2^{i+1}M$
for all $i\in\mathbb{N}$. ∎
Recall that a positive operator $T$ from a vector lattice to a normed lattice is $\sigma-Lebesgue$ if and only if $T$ is $\sigma-order-to-norm continuous$ [5, Lemma 2.1]. The same result holds also in the following more general setting.
@@
$T\in{\cal L}_{Leb}(X,Y)\Longleftrightarrow T\in{\cal L}_{otc}(X,Y)\ \ \ and\ \ \ T\in{\cal L}^{\sigma}_{Leb}(X,Y)\Longleftrightarrow T\in{\cal L}^{\sigma}_{otc}(X,Y).$
@@
As the $\sigma$-Lebesgue case is similar, we consider only the first equivalence. Since obviously ${\cal L}_{otc}(X,Y)\subseteq{\cal L}_{Leb}(X,Y)$, it suffices to show that $T\in{\cal L}_{Leb}(X,Y)$ implies $T\in{\cal L}_{otc}(X,Y)$.
On the way to a contradiction, suppose $T\in{\cal L}_{Leb}(X,Y)\setminus{\cal L}_{otc}(X,Y)$. Then, there exist a net $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$ and a neighborhood $U\in\tau(0)$ such that, for every $\alpha$, there exists $\alpha^{\prime}\geq\alpha$ with $Tx_{\alpha^{\prime}}\notin U$. Since $Y_{+}$ is normal, we may suppose that $U$ is circled and full by [4, Corollary 2.21].
Pick a net $g_{\beta}\downarrow 0$ in $X$ such that, for each $\beta$, there exists $\alpha_{\beta}$ with $x_{\alpha}\in[-g_{\beta},g_{\beta}]$ for $\alpha\geq\alpha_{\beta}$. As $T$ is Lebesgue, $Tg_{\beta}\operatorname{\xrightarrow[]{\tau}}0$. Since $T\geq 0$, $Tg_{\beta}\downarrow\geq 0$. Suppose $Tg_{\beta}\geq f\geq 0$ for some $f\in Y$. Since $\tau$ is a Hausdorff topology, it follows from [4, Theorem 2.23] that $f=0$, and hence $Tg_{\beta}\downarrow 0$. Since $T\geq 0$, we obtain $Tx_{\alpha}\in[-Tg_{\beta},Tg_{\beta}]$ for $\alpha\geq\alpha_{\beta}$. Since $Tg_{\beta}\operatorname{\xrightarrow[]{\tau}}0$, there exists $\beta_{U}$ such that $Tg_{\beta}\in U$ for all $\beta\geq\beta_{U}$. Since $U$ is circled and full, $Tg_{\beta_{U}},Tg_{\beta_{U}}]\subseteq U$ for all $\beta\geq\beta_{U}$, and hence $Tx_{\alpha}\in U$ for all $\alpha\geq\alpha_{\beta_{U}}$. It is absurd because $Tx_{\alpha^{\prime}_{\beta_{U}}}\notin U$. Thus, $T\in{\cal L}_{otc}(X,Y)$. ∎
@@
By [5, Proposition 2.2], a positive order continuous operator $T$ from a vector lattice to a normed lattice is Lebesgue if and only if $T$ is w-Lebesgue. We extend this fact to the OVS-setting as follows.
@@
Question 3.8.
Are the conclusions of Corollaries 3.5 and 3.7 hold true for arbitrary (not necessarily positive) operators?
The following lemma is a direct consequence of Definition 3.1.
@@
${\cal L}_{otc}(X,Y)\subseteq{\cal L}_{owc}(X,Y),\ \ {\cal L}^{\sigma}_{otc}(X,Y)\subseteq{\cal L}^{\sigma}_{owc}(X,Y),$
${\cal L}_{otc}(X,Y)\subseteq{\cal L}_{Leb}(X,Y)\bigcap{\cal L}^{\sigma}_{otc}(X,Y)\subseteq{\cal L}_{Leb}(X,Y)\bigcup{\cal L}^{\sigma}_{otc}(X,Y)\subseteq{\cal L}^{\sigma}_{Leb}(X,Y),\ \ \text{\rm and}$
@@
Let $X$ be an OBS with a closed generating normal cone and let $Y$ be an NS. Then ${\cal L}_{onb}(X,Y)={\cal L}_{b}(X,Y)$.
@@
The inclusion ${\cal L}_{onb}(X,Y)\subseteq{\cal L}_{b}(X,Y)$ follows from [15, Corollary 2.9]. The reverse inclusion hods due to normality of $X$. ∎
By [15, Theorem 2.8], every order-to-norm bounded operator from an OBS with closed generating cone to an NS is bounded. Accordingly to [15, Theorem 2.4], each order bounded operator from an OVS to an OVS with generating cones is ru-continuous. Combining these two results and keeping in mind that order intervals in an OBS with normal cone are bounded, we obtain the following automatic boundedness conditions of order bounded operators:
@@
    ${\cal L}_{ob}(X,Y)\subseteq{\cal L}_{b}(X,Y)$, whenever $X$ and $Y$ are OBSs with generating cones such that $X_{+}$ is closed and $Y_{+}$ is normal.
Another case of automatic boundedness of operators concerns the order continuity. Namely, as every order continuous operator from an Archimedean OVS with generating cone to an OVS is order bounded [15, Corollary 2.2], we have:
@@
    ${\cal L}_{oc}(X,Y)\subseteq{\cal L}_{b}(X,Y)$, whenever $X$ and $Y$ are OBSs with generating cones such that $X_{+}$ is closed and $Y_{+}$ is normal.
Condition $X_{+}-X_{+}=X$ is essential in both (*) and (**) since each $T:X\to X$ is order bounded and order continuous whenever $X_{+}=\{0\}$. An inspection of [15, Example 2.12 b)] tells us that each infinite dimensional Banach space $X$ has a non-closed generating Archimedean cone so that every operator $T:X\to X$ is order continuous, and hence order bounded.
@@
Combining Lemmas 3.9, 3.10, and 3.11, we obtain.
@@
Let $X$ be an ordered Banach space with a closed generating normal cone and let $Y$ be a normed space. Then ${\cal L}^{\sigma}_{wLeb}(X,Y)\subseteq{\cal L}_{b}(X,Y)$. In particular, ${\cal L}^{\sigma}_{ows}(X,Y)\subseteq{\cal L}_{b}(X,Y)$.
@@
(cf. [5, Proposition 3.2], [22, Lemma 2.3]) Each $\sigma$-w-Lebesgue, and hence each $\sigma$-order-to-weak continuous, operator from a Banach lattice to an NS is bounded.
@@
Let $X$ be an OBS with a closed generating normal cone and let $Y$ be an NS. Then ${\cal L}_{Leb}(X,Y)$, ${\cal L}^{\sigma}_{Leb}(X,Y)$, ${\cal L}_{wLeb}(X,Y)$, ${\cal L}^{\sigma}_{wLeb}(X,Y)$, ${\cal L}_{onc}(X,Y)$, ${\cal L}^{\sigma}_{onc}(X,Y)$, ${\cal L}_{owc}(X,Y)$, and ${\cal L}^{\sigma}_{owc}(X,Y)$ are norm closed subspaces of ${\cal L}_{b}(X,Y)$.
Positive order-to-norm continuous order-to-norm bounded operators between normed lattices are not necessary bounded (see [5, Remark 1.1. c)]). From the other hand, even a positive $\sigma$-Lebesgue rank one contraction on an AM-space need not to be w-Lebesgue [5, Example 3.1].